# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 9.3 Eigenvalue crossing
Figure 52: (a) A small change in the bath temperature $T_{b}$ drastically changes the approach to equilibrium direction in the presence of a level crossing like the one observed in the four-wells setup shown in panel (b), representing the spectrum of an $N=8$ Ising antiferromagnet at zero magnetic field with Glauber dynamics [79]. Highlighted in blue and red are the first two dominant eigenvalues, which are relevant for the dynamics. They cross at $T^{*}(N=8)\sim 5.45$. (c) Eigenvalue difference $\delta\lambda_{23}$ as a function of the reduced temperature $t=(T-T^{*})/T^{*}$, exhibiting an excellent finite-size collapse $\propto|t|^{1}$ around the crossing temperature. Inset: the distance from the asymptotic crossing temperature $T^{*}_{\infty}=2/\mathrm{arctanh}(1/3)$ decays quadratically. Figures adapted from [82].
@@
Eigenvalue crossing can be viewed as a phase transition in the relaxation dynamics, which occurs in the long-time limit $t\to\infty$, in analogy with an equilibrium phase transition, which can be observed in the thermodynamic limit $N\to\infty$. From the equation regulating the last stages of the evolution in a thermal quench (Eq. (14)), one sees how the timescale is regulated by an exponential decay with characteristic time $\propto 1/|\lambda_{2}|$. To highlight this, we can formally rearrange the dynamics equation, obtaining
@@
$\vec{v}_{2}=\begin{cases}\vec{v}^{-}_{2}&T_{b}<T^{*}\\ a_{2}^{-}\vec{v}^{-}_{2}+a_{2}^{+}\vec{v}^{+}_{2}&T_{b}=T^{*}\\ \vec{v}^{+}_{2}&T_{b}>T^{*}\\ \end{cases}$ (155)
@@
While eigenvalue crossing can be observed in very small systems (in Ref. [82] the example of a 4-state Markovian system is presented), it can also be observed in complex many-body interacting systems as Ising antiferromagnets in zero external magnetic fields $H=0$. For a 1D periodic chain of $N$ spins with Glauber dynamics and $H=0$, there is an exact solution of the spectrum $\lambda_{k}(T_{b})$ (with $k=1,\dots,2^{N}$) [80]. This solution allows visualization (see Fig. 52b for the case $N=8$) of how the spectrum evolves as a function of the bath temperature $T_{b}$, highlighting a high number of crossing among the eigenvalues. We highlighted in red and blue the relevant eigenvalues $\lambda_{2}^{\pm}$, which are those that are non-orthogonal to all Boltzmann distributions. The higher eigenvalues are orthogonal to all Boltzmann distributions. Thus, equilibrium initial conditions have zero projection on these and are hence irrelevant to our discussion. Their analytic expressions for finite $N$ are given by
$\lambda_{2}^{-}$ $=-2+2\tanh(2J/T_{b})\cos\left(\frac{N-1}{N}\pi\right),$ (156)
$\lambda_{2}^{+}$ $=-1+\tanh(2J/T_{b}),$ (157)
@@
The presence of the above eigenvalue crossing at $H=0$ is directly connected [70] with the emergence of a Mpmeba effect in antiferromagnets for $H\ll 1$ above a certain temperature (see Fig. 11b). Indeed, such a crossing dramatically changes the coefficient of the second eigenvector in the initial condition and hence affects its monotonicity through which the effect is defined. The crossing becomes progressively sharper in the anti-ferromagnetic system as the magnetic field $H$ goes to zero. This feature was observed transversely across different Ising antiferromagnets systems and different dynamics, for which the effect does not seem to appear for any value of the two external parameters—bath temperature $T_{b}$ and magnetic field $H$—but rather in a specific domain (e.g., compare the phase diagrams for a mean-field model [96] and a 1D model subject to boundary coupling [70]).
@@
Figure 53: Projection of the slowest relaxation vector $\vec{v}_{2}$ on the magnetization (a) and staggered magnetization (b) for an $N=8$ one-dimensional antiferromagnetic spin chain (see the main text for the specifics of the model). In panel (c) we highlight the areas in which the projection of $\vec{v}_{2}$ is larger along the magnetization (red area) and staggered magnetization (blue area). (d,e) Equilibrium loci showing the values of magnetization and staggered magnetization of the chain in equilibrium at different temperatures $T_{b}$ below ($H=1.95$) and above ($H=2.05$) the critical value of the magnetization that determines the phase transition in the groundstate of the Ising chain. This figure is reprinted from [82].
@@
To better understand the connection between these two effects, an important observation is related to the fact that the eigenvectors associated with the crossing eigenvalues carry a physical meaning: $\vec{v}_{2}$ is parallel to the staggered magnetization ($\vec{m}_{\rm s}$) before the crossing (where no Mpemba effect is observed) while it is parallel to the magnetization vector ($\vec{m}_{\rm s}$) after the crossing (where strong inverse Mpemba effect can be observed) [82]. For small magnetic fields (such that the antiferromagnetic interaction dominates), the magnetization of an Ising antiferromagnet at equilibrium is a nonmonotonic observable with respect to the temperature, while the staggered magnetization at equilibrium, on the other hand, is monotonous with respect to the temperature (see Fig. 53d). In Figs. (53)a and (53)b the projection of $\vec{v}_{2}$ on $\vec{m}$ and $\vec{m}_{\rm s}$ respectively are plotted, highlighting the magnitude of the projections, while in Fig. 53c we color in red the area in which $|\vec{v}_{2}\cdot\vec{m}|>|\vec{v}_{2}\cdot\vec{m}_{\rm s}|$ and blue the one in which $|\vec{v}_{2}\cdot\vec{m}|>|\vec{v}_{2}\cdot\vec{m}_{\rm s}|$. This explains geometrically why, having the slowest relaxation occurring along the magnetization for $H<2$, is a prerequisite in order to observe a strong Mpemba effect, and consequently why there is no effect before the crossing.
@@
For large magnetic fields (such that the field overcomes the antiferromagnetic character), the opposite holds: the staggered magnetization becomes nonmonotonic while the magnetization is now monotonic (Fig. 53e). Indeed, strong Mpemba effect can be observed in this case (see Refs. [70, 5]) only in the blue area above $H>2$. This demonstrates clearly that the Mpemba effect is a phenomenon intrinsically related to the dynamics, as discussed in Sec. 2.1.9: indeed, the geometric shortcut that is enabled by the nonmonotonicity of an observable as in Fig. 53 (see also Fig. 13 in the case of a weakly coupled system) is not a two-way road, as the spectrum of the rate matrix explicitly depends on $T_{b}$ and can therefore be dramatically different in the two cases.
@@ 9.4 Relaxation speedup
The following sections discuss two methods for accelerating relaxation processes. These methods utilize the time evolution of physical observables and the complex internal structures of physical systems near a phase transition. The first approach, outlined in Sec. 9.4.1, shifts the focus from the evolution of the probability vector (or density) to the evolution of physical observables. This exploits the relationship between two formulations of the dynamics: one in which vectors carry the time dependence and the other in which time dependence is expressed through observables. A quantum mechanical analogy of this is switching from a formulation with time-dependent states to one with time-dependent observables. The second approach, presented in Sec. 9.4.2, leverages the complex internal structure near a phase transition to reduce the time required to reach a target state. We illustrate both approaches using 1D and 2D models of Ising spins with antiferromagnetic interactions.
@@ 9.4.1 Speedup equilibrium properties
Figure 54: An observable whose equilibrium expected value exhibits a nonmonotonic dependence on temperature can be utilized to optimize its value near the maximum. Two illustrative examples are presented in the panels. Panel (a), on the left, depicts the Mpemba effect, while the right side illustrates the speedup achieved by briefly heating the system before cooling it. The red lines represent the hotter system, while the blue lines indicate the colder system. For a 1D Ising antiferromagnetic model, panel (b) left reproduces the Mpemba effect in the change of sing of $\Delta_{t}\langle\varepsilon\rangle=(\langle\varepsilon(t)\rangle^{h}-\langle\varepsilon(t)\rangle^{c})/N$ (with $\varepsilon$ defined in Eq. (44) and $c$ and $h$ indicates cold and hot respectively). Panel (b) right shows the supercooling produced by the preheating strategy for $\delta_{t}\langle m_{s}^{2}\rangle=(\langle m_{s}^{2}(t)\rangle-\langle m_{s}^{2}(t)\rangle^{b})/N$ the squared staggered magnetization, where the superindex $b$ indicates the bath. Figures are reprinted from  [283].
As we have seen throughout the review, the various Mpemba effects can reduce the time needed to cool down (or heat up) a system, as has been shown, for example, in Markovian systems that satisfy detailed balance (Sec. 2). The analysis of the speedup relies on the diagonalization of the transition-rate matrix of the Master Eq. (6) and finding cases in which the probability vector $\vec{p}(t)$ has no contribution from the eigenvector $\vec{v}_{2}$ (which describes the slowest evolving direction in probability space), thus forcing an exponential speedup towards equilibrium. In  [283], the authors propose studying the weak version of the Master Eq. (6). The idea, similar in spirit (though not in detail) to the numerical method introduced in Sec. 7.1.3, is to focus on how different observables evolve in time instead of focusing on the probability vector itself. Related to the eigenbasis $(\vec{\pi},\vec{v}_{2},\vec{v}_{3},\ldots)$, there is a dual basis of observables $\left(\boldsymbol{1},\mathcal{O}^{\mathrm{b}}_{2},\mathcal{O}^{\mathrm{b}}_{3},\ldots\right)$. Analogously, this allows one to write the evolution of the expected value of an observable $\mathcal{A}$ as
@@
Note that Eq. (158) is akin to Eq. (12) but using an observable instead of a probability—much like the differences between the Schrödinger and Heisenberg representations in quantum mechanics [284]. An observable shows an exponential speedup if it initially has no contribution from $\mathcal{O}^{\mathrm{b}}_{2}$, as that is the slowest evolving observable mode. However, the observables $\mathcal{O}^{\mathrm{b}}_{k}$ are mere mathematical constructs. That is, they might not represent macroscopic physical observables in the thermodynamic limit. To overcome this shortcoming, the authors propose general requirements for the system based on timescale separation and nonmonotonic temperature evolution of an important state function. These features are quite generic near a first-order phase transition.
@@
With this in mind, the authors explain the appearance of several paradoxical behaviors near that maximum (cf. Fig. 54). This also serves the purpose of finding effective strategies to devise such phenomena in mesoscopic systems. In particular, the model system studied in [283] is the antiferromagnetic 1D Ising model given in Eq. (44). In addition, the equilibrium observable used by the authors is the squared staggered magnetization (described in Section 3 as $m_{s}=\sum_{k=1}^{N}(-1)^{k}\sigma_{k}$). The three found anomalous effects are (i) the Mpemba effect, (ii) the Preheating strategy, and (iii) two asymmetries in heating and cooling processes, namely, heating faster than cooling and vice versa, as discussed in section 9.2.In Fig. 54, two of these anomalous effects are shown.
@@ 9.4.2 Internal structure speedup
Often, a system has a complex internal structure near a (first- or second-order) phase transition. In Ref. [285], the authors show that the slow growth of magnetic domains helps to shorten the time it takes the system to reach a desired state. To this end, they consider a 2D Ising model without a magnetic field, described by the Hamiltonian $\mathcal{H}=-J\sum_{{\rm n.n.}\,x,y}\sigma_{x}\sigma_{y}$, where spin $\sigma_{x}$ has position $x$ in a lattice $L\times L$. To demonstrate their idea, the authors study the 2D Ising Model (with $L=4096$) using extensively two simulation methods, the Metropolis and the Heat Bath algorithms. This model has no conserved magnitudes and belongs to the model A universality class, where the dynamics has no conserved quantities [286]. A second-order phase transition between the ferromagnetic and paramagnetic phase appears at $T=T_{\rm c}$. It is well known that in the ferromagnetic phase and excluding fast initial relaxations, the energy $E(t)$ and the correlation length $\xi(t)$ are related because the excess energy is located at the boundaries of the magnetic domains.
Figure 55: Example of anomalous relaxation in a 2D Ising model, using a two-step protocol for $\kappa_{\rm start}=0.46$ and $\kappa_{\rm end}=0.4378$. (a) Change of sign of $\mathcal{\xi}$ and an example of $\xi^{*}_{\rm start}$. (b) Examples of the speedup using a two-step protocol. (c) Evolution of $\xi(t)$ in the one-step protocol. Note that the largest timescale exists as a finite-size effect in the ferromagnetic phase. Indeed, when $\kappa>\kappa_{\rm c}$ and keeping leading terms, domains grow as $\xi(t)\approx\sqrt{t}$ until $\xi\approx L$. (d) Equilibration time $t_{\rm eq}$ as a function of $\xi_{\rm start}$. Note that the equilibration time $t_{\rm eq}$ here is defined as the time such that $E_{\rm total}$ differs from $E_{\rm eq}$ by less than $0.1\%$ for any $t_{\rm total}>t_{\rm eq}$. MET and HB indicate the use of Metropolis and Heat Bath algorithms, respectively. The figures are adapted from [285].
@@
where $\tau_{k}$ is the largest relaxation timescale. Here, $\mathcal{A}(\xi)$ is an amplitude that depends only on the coherence length $\xi$. Interestingly, $\mathcal{A}(\xi)$ change its sign as the initial coherence length grows, and $\mathcal{A}(\xi^{*}_{\rm start})=0$ at some point; see Fig. 55a. Given that $\xi(t)$ grows faster in the ferromagnetic phase (see Fig. 55a), it is likely that $E(t)$ could equilibrate faster in the paramagnetic phase through an excursion to the ferromagnetic one. In addition, an exponential speedup is possible if the optimal $\xi^{*}_{\rm start}$ is found. Two thermal protocols are designed, with one and two steps, to study that phenomenon. In the one-step protocol, a disordered spin configuration, corresponding to infinite temperature, is put in contact with a thermal bath at inverse-temperature $\kappa=J/T$; see Fig. 55c. The equilibrium value of $\xi$ is reached only in the paramagnetic phase. In the two-step protocol, a fully disordered spin configuration is initially placed at an inverse temperature in the ferromagnetic phase, $\kappa_{\rm start}>\kappa_{\mathrm{c}}$, where it evolves until $\xi(t)$ reaches a target value $\xi_{\rm start}(t)$. Just at that point, the bath is changed to a new $\kappa_{\rm end}<\kappa_{\mathrm{c}}$. Thus, the second change must be done when $\mathcal{A}(\xi^{*}_{\rm start})=0$, $\xi^{*}_{\rm start}$ can be estimated numerically. From Fig. 55a and Fig. 55b, the speedup due to the choice of $\xi^{*}_{\rm start}=78$ can be clearly observed. Interestingly, for $\xi^{*}_{\rm start}=7.92$, the Kovacs effect is observed; cf. Sec. 9.1.

### v3

@@ 9.3 Eigenvalue crossing
Figure 52: (a) A small change in the bath temperature $T_{b}$ drastically changes the approach to equilibrium direction in the presence of a level crossing like the one observed in the four-wells setup shown in panel (b), representing the spectrum of an $N=8$ Ising antiferromagnet at zero magnetic field with Glauber dynamics [79]. Highlighted in blue and red are the first two dominant eigenvalues, which are relevant for the dynamics. They cross at $T^{*}(N=8)\sim 5.45$. (c) Eigenvalue difference $\delta\lambda_{23}$ as a function of the reduced temperature $t=(T-T^{*})/T^{*}$, exhibiting an excellent finite-size collapse $\propto|t|^{1}$ around the crossing temperature. Inset: the distance from the asymptotic crossing temperature $T^{*}_{\infty}=2/\mathrm{arctanh}(1/3)$ decays quadratically. Source: Adapted with permission from [82].
@@
Eigenvalue crossing can be viewed as a phase transition in the relaxation dynamics, which occurs in the long-time limit $t\to\infty$, in analogy with an equilibrium phase transition, which can be observed in the thermodynamic limit $N\to\infty$. From the equation regulating the last stages of the evolution in a thermal quench, Eq. (14), one sees how the timescale is regulated by an exponential decay with characteristic time $\propto 1/|\lambda_{2}|$. To highlight this, we can formally rearrange the dynamics equation, obtaining
@@
$\vec{v}_{2}=\begin{cases}\vec{v}^{-}_{2}&T_{b}<T^{*}\\ a_{2}^{-}\vec{v}^{-}_{2}+a_{2}^{+}\vec{v}^{+}_{2}&T_{b}=T^{*}\\ \vec{v}^{+}_{2},&T_{b}>T^{*}\\ \end{cases}$ (155)
@@
While eigenvalue crossing can be observed in very small systems (e.g., in Ref. [82], where the example of a 4-state Markovian system is presented), it can also be observed in complex many-body interacting systems as Ising antiferromagnets in zero external magnetic fields $H=0$. For a 1D periodic chain of $N$ spins with Glauber dynamics and $H=0$, there is an exact solution of the spectrum $\lambda_{k}(T_{b})$ (with $k=1,\dots,2^{N}$) [80]. This solution allows visualization (see Fig. 52b for the case $N=8$) of how the spectrum evolves as a function of the bath temperature $T_{b}$, highlighting a high number of crossing among the eigenvalues. We highlighted in red and blue the relevant eigenvalues $\lambda_{2}^{\pm}$, which are those that are non-orthogonal to all Boltzmann distributions. The higher eigenvalues are orthogonal to all Boltzmann distributions. Thus, equilibrium initial conditions have zero projection on these and are hence irrelevant to our discussion. Their analytic expressions for finite $N$ are given by
$\lambda_{2}^{-}$ $=-2+2\tanh(2J/k_{B}T_{b})\cos\left(\frac{N-1}{N}\pi\right),$ (156)
$\lambda_{2}^{+}$ $=-1+\tanh(2J/k_{B}T_{b}),$ (157)
@@
The presence of the above eigenvalue crossing at $H=0$ is directly connected [70] with the emergence of a Mpmeba effect in antiferromagnets for $H\ll 1$ above a certain temperature (see Fig. 11b). Indeed, such a crossing dramatically changes the coefficient of the second eigenvector in the initial condition and hence affects its monotonicity through which the effect is defined. The crossing becomes progressively sharper in the antiferromagnetic system as the magnetic field $H$ goes to zero. This feature was observed transversely across different Ising antiferromagnets systems and different dynamics, for which the effect does not seem to appear for any value of the two external parameters—bath temperature $T_{b}$ and magnetic field $H$—but rather in a specific domain (e.g., compare the phase diagrams for a mean-field model [96] and a 1D model subject to boundary coupling [70]).
@@
Figure 53: Projection of the slowest relaxation vector $\vec{v}_{2}$ on the magnetization (a) and staggered magnetization (b) for an $N=8$ 1D antiferromagnetic spin chain; see the main text for the specifics of the model. (c) Highlight of the areas in which the projection of $\vec{v}_{2}$ is larger along the magnetization (red area) and staggered magnetization (blue area). (d,e) Equilibrium loci showing the values of magnetization and staggered magnetization of the chain in equilibrium at different temperatures $T_{b}$ below ($H=1.95$) and above ($H=2.05$) the critical value of the magnetization that determines the phase transition in the groundstate of the Ising chain. Source: Reprinted with permission from [82].
@@
To better understand the connection between these two effects, an important observation is related to the fact that the eigenvectors associated with the crossing eigenvalues carry a physical meaning: $\vec{v}_{2}$ is parallel to the staggered magnetization ($\vec{m}_{\rm s}$) before the crossing, where no Mpemba effect is observed, while it is parallel to the magnetization vector ($\vec{m}_{\rm s}$) after the crossing, where a strong inverse Mpemba effect can be observed [82]. For small magnetic fields (such that the antiferromagnetic interaction dominates), the magnetization of an Ising antiferromagnet at equilibrium is a nonmonotonic observable with respect to the temperature, while the staggered magnetization at equilibrium, on the other hand, is monotonic with respect to the temperature; see Fig. 53d. In Figs. 53a and 53b, the projection of $\vec{v}_{2}$ on $\vec{m}$ and $\vec{m}_{\rm s}$ respectively are plotted, highlighting the magnitude of the projections, while in Fig. 53c we color in red the area in which $|\vec{v}_{2}\cdot\vec{m}|>|\vec{v}_{2}\cdot\vec{m}_{\rm s}|$ and blue the one in which $|\vec{v}_{2}\cdot\vec{m}|>|\vec{v}_{2}\cdot\vec{m}_{\rm s}|$. This explains geometrically why, having the slowest relaxation occurring along the magnetization for $H<2$, is a prerequisite in order to observe a strong Mpemba effect, and consequently why there is no effect before the crossing.
@@
For magnetic fields that are large enough to overcome the antiferromagnetic ordering, the opposite holds: the staggered magnetization becomes nonmonotonic while the magnetization is now monotonic; see Fig. 53e. Indeed, strong Mpemba effect can be observed in this case only in the blue area above $H>2$; see Refs. [70, 5]. This demonstrates clearly that the Mpemba effect is a phenomenon intrinsically related to the dynamics, as discussed in Sec. 2.1.9: indeed, the geometric shortcut that is enabled by the nonmonotonicity of an observable as in Fig. 53 (see also Fig. 13 in the case of a weakly coupled system) is not a two-way road, as the spectrum of the rate matrix explicitly depends on $T_{b}$ and can therefore be dramatically different in the two cases.
@@ 9.4 Relaxation speedup
The following sections discuss two methods for accelerating relaxation processes. These methods use the time evolution of physical observables and the complex internal structures of physical systems near a phase transition. The first approach, outlined in Sec. 9.4.1, shifts the focus from the evolution of the probability vector (or density) to the evolution of physical observables. This exploits the relationship between two formulations of the dynamics, one in which vectors carry the time dependence and the other in which time dependence is expressed through observables. A quantum-mechanical analogy of this is switching from a formulation with time-dependent states to one with time-dependent observables. The second approach, presented in Sec. 9.4.2, leverages the complex internal structure near a phase transition to reduce the time required to reach a target state. We illustrate both approaches using 1D and 2D models of Ising spins with antiferromagnetic interactions.
@@ 9.4.1 Speedup equilibrium properties
Figure 54: An observable whose equilibrium expected value exhibits a nonmonotonic dependence on temperature can be utilized to optimize its value near the maximum. Two illustrative examples are presented in the panels. Panel (a), on the left, depicts the Mpemba effect, while the right side illustrates the speedup achieved by briefly heating the system before cooling it. The red lines represent the hotter system, while the blue lines indicate the colder system. For a 1D Ising antiferromagnetic model, panel (b) left reproduces the Mpemba effect in the change of sing of $\Delta_{t}\langle\varepsilon\rangle=(\langle\varepsilon(t)\rangle^{h}-\langle\varepsilon(t)\rangle^{c})/N$ (with $\varepsilon$ defined in Eq. (44) and $c$ and $h$ indicates cold and hot respectively). Panel (b) right shows the supercooling produced by the preheating strategy for $\delta_{t}\langle m_{s}^{2}\rangle=(\langle m_{s}^{2}(t)\rangle-\langle m_{s}^{2}(t)\rangle^{b})/N$ the squared staggered magnetization, where the superindex $b$ indicates the bath. Source: Reprinted with permission from [296].
As we have seen throughout the review, the various Mpemba effects can reduce the time needed to cool down (or heat up) a system, as has been shown, for example, in Markovian systems that satisfy detailed balance (Sec. 2). The analysis of the speed up relies on the diagonalization of the transition-rate matrix of the master Eq. (6) and finding cases in which the probability vector $\vec{p}(t)$ has no contribution from the eigenvector $\vec{v}_{2}$ (which describes the slowest evolving direction in probability space), thus forcing an exponential speed up towards equilibrium. In  [296], the authors propose studying the weak version of the master Eq. (6). The idea is similar in spirit, although not in detail, to the numerical method introduced in Sec. 7.1.3. We focus on how different observables evolve in time instead of focusing on the probability vector itself. Related to the eigenbasis $(\vec{\pi},\vec{v}_{2},\vec{v}_{3},\ldots)$, there is a dual basis of observables $\left(\boldsymbol{1},\mathcal{O}^{\mathrm{b}}_{2},\mathcal{O}^{\mathrm{b}}_{3},\ldots\right)$. Analogously, this allows one to write the evolution of the expected value of an observable $\mathcal{A}$ as
@@
Note that Eq. (158) is akin to Eq. (12) but using an observable instead of a probability—much like the differences between the Schrödinger and Heisenberg representations in quantum mechanics [297]. An observable shows an exponential speedup if it initially has no contribution from $\mathcal{O}^{\mathrm{b}}_{2}$, as that is the slowest evolving observable mode. However, the observables $\mathcal{O}^{\mathrm{b}}_{k}$ are mere mathematical constructs. That is, they might not represent macroscopic physical observables in the thermodynamic limit. To overcome this shortcoming, the authors propose general requirements for the system based on timescale separation and nonmonotonic temperature evolution of an important state function. These features are generic near a first-order phase transition.
@@
With this in mind, the authors explain the appearance of several paradoxical behaviors near that maximum; cf. Fig. 54. This also serves the purpose of finding effective strategies to devise such phenomena in mesoscopic systems. In particular, the model system studied in [296] is the antiferromagnetic 1D Ising model given in Eq. (44). In addition, the equilibrium observable used by the authors is the squared staggered magnetization (described in Sec. 3 as $m_{s}=\sum_{k=1}^{N}(-1)^{k}\sigma_{k}$). The three found anomalous effects are (i) the Mpemba effect, (ii) the Preheating strategy, and (iii) two asymmetries in heating and cooling processes, namely, heating faster than cooling and vice versa, as discussed in Sec. 9.2. In Fig. 54 two of these anomalous effects are shown.
@@ 9.4.2 Internal structure speedup
Often, a system has a complex internal structure near a (first- or second-order) phase transition. In [298], the authors show that the slow growth of magnetic domains helps to shorten the time it takes the system to reach a desired state. To this end, they consider a 2D Ising model without a magnetic field, described by the Hamiltonian
$\mathcal{H}=-J\displaystyle\sum_{{\rm n.n.}\,x,y}\sigma_{x}\sigma_{y},$ (161)
where spin $\sigma_{x}$ has position $x$ in a lattice $L\times L$. To demonstrate their idea, the authors study the 2D Ising Model (with $L=4096$) using extensively two simulation methods, the Metropolis and the heat-bath algorithms. This model has no conserved magnitudes and belongs to the model A universality class, where the dynamics has no conserved quantities [299]. A second-order phase transition between the ferromagnetic and paramagnetic phase appears at $T=T_{\rm c}$. It is well known that in the ferromagnetic phase and excluding fast initial relaxations, the energy $E(t)$ and the correlation length $\xi(t)$ are related because the excess energy is located at the boundaries of the magnetic domains.
Figure 55: Example of anomalous relaxation in a 2D Ising model, using a two-step protocol for $\kappa_{\rm start}=0.46$ and $\kappa_{\rm end}=0.4378$. (a) Change of sign of $\mathcal{\xi}$ and an example of $\xi^{*}_{\rm start}$. (b) Examples of the speedup using a two-step protocol. (c) Evolution of $\xi(t)$ in the one-step protocol. Note that the largest timescale exists as a finite-size effect in the ferromagnetic phase. Indeed, when $\kappa>\kappa_{\rm c}$ and keeping leading terms, domains grow as $\xi(t)\approx\sqrt{t}$ until $\xi\approx L$. (d) Equilibration time $t_{\rm eq}$ as a function of $\xi_{\rm start}$. Note that the equilibration time $t_{\rm eq}$ here is defined as the time such that $E_{\rm total}$ differs from $E_{\rm eq}$ by less than $0.1\%$ for any $t_{\rm total}>t_{\rm eq}$. MET and HB indicate the use of Metropolis and Heat Bath algorithms, respectively. Source: Adapted with permission from [298].
@@
where $\tau_{k}$ is the largest relaxation timescale. Here, $\mathcal{A}(\xi)$ is an amplitude that depends only on the coherence length $\xi$. Interestingly, $\mathcal{A}(\xi)$ change its sign as the initial coherence length grows, and $\mathcal{A}(\xi^{*}_{\rm start})=0$ at some point; see Fig. 55a. Given that $\xi(t)$ grows faster in the ferromagnetic phase (see Fig. 55a), it is likely that $E(t)$ could equilibrate faster in the paramagnetic phase through an excursion to the ferromagnetic one. In addition, an exponential speedup is possible if the optimal $\xi^{*}_{\rm start}$ is found. Two thermal protocols are designed, with one and two steps, to study that phenomenon. In the one-step protocol, a disordered spin configuration, corresponding to infinite temperature, is put in contact with a thermal bath at inverse-temperature $\kappa=J/k_{B}T$; see Fig. 55c. The equilibrium value of $\xi$ is reached only in the paramagnetic phase. In the two-step protocol, a fully disordered spin configuration is initially placed at an inverse temperature in the ferromagnetic phase, $\kappa_{\rm start}>\kappa_{\mathrm{c}}$, where it evolves until $\xi(t)$ reaches a target value $\xi_{\rm start}(t)$. Just at that point, the bath is changed to a new $\kappa_{\rm end}<\kappa_{\mathrm{c}}$. Thus, the second change must be done when $\mathcal{A}(\xi^{*}_{\rm start})=0$, where $\xi^{*}_{\rm start}$ can be estimated numerically. From Fig. 55a and Fig. 55b, the speedup due to the choice of $\xi^{*}_{\rm start}=78$ can be clearly observed. Interestingly, for $\xi^{*}_{\rm start}=7.92$, the Kovacs effect is observed; cf. Sec. 9.1.